OpenStudy (anonymous):

Find the domain and range of f(x)=(2x+3)/(x-5) using f(x) and the inverse of f(x).

OpenStudy (anonymous):

domain: \[x=\mathbb{R}\]\[x \neq5\]

OpenStudy (anonymous):

and whats thee range.

OpenStudy (anonymous):

um. i don't think there's a limit on the range but i'm not entirely sure.

hero (hero):

if there's a limit on x, there's a limit on y

OpenStudy (anonymous):

So, domain =\[(-\infty,5)\cup(5,\infty)\] and range =\[(-\infty,\infty)\]

hero (hero):

There has to be a limit on y if there's a limit on x

hero (hero):

you can't evaluate f(5)

OpenStudy (anonymous):

no there doesn't. it just means that y isn't undefined

hero (hero):

I'll find out what y can't equal. All I have to do is just find the inverse

hero (hero):

since f(x) has an inverse, that proves that there's a limit on y

OpenStudy (anonymous):

just look at the sine function, there's a limit on the range but not on the x

OpenStudy (anonymous):

so what's thee range?

hero (hero):

y cannot equal 2

hero (hero):

liamarie, you're saying it backwards. With the sine function x has no limit. But what I said earlier is that if there is a limit on x, there's a limit on y. It doesn't work the other way around.

OpenStudy (anonymous):

Can I ask how you got 2?

hero (hero):

I found the inverse of f(x)

hero (hero):

f^-1(x) = (5x+3)/(x-2)

OpenStudy (anonymous):

hm. i stand corrected.

OpenStudy (anonymous):

wouldnt it be, (5x-3)/(x+2)

hero (hero):

No. I don't know how you got that. But no

OpenStudy (anonymous):

ahah okay, thanks.

hero (hero):

We can prove that f(x) cannot equal 2 by setting f(x) = 2 and solving for x

hero (hero):

Do that and see what happens.